What is Sq Rt(x-1) SQUARED? what I am asking is that the problem takes the square root of (x-1) and then squares it. Is it x-1 or x+1.
If you could answer the whole problem that would be appreciated, but it is a calculus question, but I am stuck on one of the algebra parts.
Thanks.

Answer by stanbon(75887)

For x>=1, [sqrt(x-1)]^2 = x-1
--------
For all x, sqrt[(x-1)^2]= |x-1|
---------
Cheers,
Stan H.